# Tree-level 11D supergravity amplitudes from the pure spinor worldline

### v1

@@ 3.2. Simplification
$\eqalignno{{\cal A}_{4}^{s}+{\cal A}_{4}^{t}&={6\over s_{12}}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{b}_{12}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+{6\over s_{13}}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{b}_{13}\phi_{2}^{a}U_{4}^{(3)}\right\rangle\cr+&12\left\langle(\lambda\gamma_{ab}\phi_{1})\phi^{a}_{2}\phi_{3}^{b}U_{4}^{(3)}\right\rangle+12\left\langle(\lambda\gamma_{ab}\phi_{2})\phi_{3}^{a}\phi_{1}^{b}U_{4}^{(3)}\right\rangle+12\left\langle(\lambda\gamma_{ab}\phi_{3})\phi_{1}^{a}\phi_{2}^{b}U_{4}^{(3)}\right\rangle\cr+&3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{1}^{a}h^{b}_{(2,a}\phi_{3)}^{a}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{1}^{a}h^{b}_{(2,\alpha}\phi_{3)}^{\alpha}U_{4}^{(3)}\right\rangle\cr+&3\left\langle C_{1a}\Big{(}\phi_{(2}^{c}\partial_{c}\phi^{a}_{3)}+\phi_{(2}^{\delta}D_{\delta}\phi^{a}_{3)}-\Omega_{(2}^{\alpha}h^{a}_{3),\alpha}\Big{)}U_{4}^{(3)}\right\rangle&\cr-&3\left\langle C_{1\alpha}\Big{(}\phi_{(2}^{c}\partial_{c}\phi^{\alpha}_{3)}+\phi_{(2}^{\delta}D_{\delta}\phi^{\alpha}_{3)}-\Omega_{(2}^{\beta}h^{\alpha}_{3),\beta}\Big{)}U_{4}^{(3)}\right\rangle.&({3.19})}$
@@
$\eqalignno{{\cal A}^{u}_{4}=&\left\langle\left\{U_{1}^{(3)},V_{23}^{(1)}\right\}U_{4}^{(3)}\right\rangle\cr=&\ {6\over s_{23}}\Big{(}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{b}_{1}\tilde{\phi}_{23}^{a}U_{4}^{(3)}\right\rangle+3\left\langle C_{1,a}(\{Q,\tilde{\phi}_{23}^{a}\}-(\lambda\gamma^{a}\tilde{\phi}_{23}))U_{4}^{(3)}\right\rangle-3\left\langle C_{1,\alpha}([Q,\tilde{\phi}_{23}^{\alpha}]-\tilde{\Omega}^{\alpha}_{3})U_{4}^{(3)}\right\rangle\Big{)}\cr=&\ {6\over s_{23}}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{b}_{1}\phi_{23}^{a}U_{4}^{(3)}\right\rangle-3\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}h^{b}_{(2,\alpha}\phi^{\alpha}_{3)}U_{4}^{(3)}\right\rangle-3\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}h^{b}_{(2,c}\phi^{c}_{3)}U_{4}^{(3)}\right\rangle\cr&-3\left\langle C_{1,a}\big{(}\phi_{(2}^{c}\partial_{c}\phi_{3)}^{a}+\phi_{(2}^{\delta}D_{\delta}\phi_{3)}^{a}-\Omega_{(2}^{\alpha}h^{a}_{3),\alpha}\big{)}U_{4}^{(3)}\right\rangle&\cr&+3\left\langle C_{1,\alpha}\big{(}\phi_{(2}^{c}\partial_{c}\phi_{3)}^{\alpha}+\phi_{(2}^{\delta}D_{\delta}\phi_{3)}^{\alpha}-\Omega_{(2}^{\beta}h^{\alpha}_{3),\beta}\big{)}U_{4}^{(3)}\right\rangle.&({3.20})}$
After plugging eqns. (3.19) and (3.20) into eqn. (3.15), one finds the simple compact formula for the 4-point function in 11D
$\eqalignno{{\cal A}_{4}=&{\cal A}_{4}^{s}+{\cal A}_{4}^{t}+{\cal A}_{4}^{u}&\cr=&\ 6\Bigg{[}\ {1\over s_{12}}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{b}_{12}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+{1\over s_{13}}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{b}_{13}\phi_{2}^{a}U_{4}^{(3)}\right\rangle+{1\over s_{23}}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{b}_{23}\phi_{1}^{a}U_{4}^{(3)}\right\rangle\cr&\ \ \ \ \ \ \ +2\left\langle(\lambda\gamma_{ab}\phi_{1})\phi^{b}_{2}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+2\left\langle(\lambda\gamma_{ab}\phi_{3})\phi^{b}_{1}\phi_{2}^{a}U_{4}^{(3)}\right\rangle+2\left\langle(\lambda\gamma_{ab}\phi_{2})\phi^{b}_{3}\phi_{1}^{a}U_{4}^{(3)}\right\rangle\Bigg{]}.&\cr&&({3.21})}$
@@ 3.4. BRST invariance and full permutation symmetry
$\eqalignno{&{\cal A}_{4}=4\left\langle V^{(3)}_{12}\{Q,V^{(3)}_{34}\}\right\rangle+4\left\langle V^{(3)}_{13}\{Q,V^{(3)}_{24}\}\right\rangle+4\left\langle V^{(3)}_{23}\{Q,V^{(3)}_{14}\}\right\rangle,&({3.22})}$
@@
$\eqalignno{\{Q,A_{4}\}=&12\left\langle(\lambda\gamma^{[ab}\lambda)(\lambda\gamma^{c]d}\lambda)\phi_{1,a}\phi_{2,b}\phi_{3,c}\phi_{4,d}\right\rangle.&({3.23})}$
@@
$\eqalignno{{\cal A}(1,2,3,4)=&4\left\langle V^{(3)}_{12}\{Q,V^{(3)}_{34}\}\right\rangle+4\left\langle V^{(3)}_{13}\{Q,V^{(3)}_{24}\}\right\rangle+4\left\langle V^{(3)}_{23}\{Q,V^{(3)}_{14}\}\right\rangle\cr=&4\left\langle\{Q,V^{(3)}_{21}\}V^{(3)}_{34}\right\rangle+4\left\langle V^{(3)}_{23}\{Q,V^{(3)}_{14}\}\right\rangle+4\left\langle V^{(3)}_{13}\{Q,V^{(3)}_{24}\}\right\rangle={\cal A}(2,1,3,4),&\cr&&({3.24})}$
@@
$\eqalignno{{\cal A}(1,2,3,4)=&4\left\langle V^{(3)}_{12}\{Q,V^{(3)}_{34}\}\right\rangle+4\left\langle V^{(3)}_{13}\{Q,V^{(3)}_{24}\}\right\rangle+4\left\langle V^{(3)}_{23}\{Q,V^{(3)}_{14}\}\right\rangle\cr=&4\left\langle\{Q,V^{(3)}_{21}\}V^{(3)}_{43}\right\rangle+4\left\langle\{Q,V^{(3)}_{31}\}V^{(3)}_{42}\right\rangle+4\left\langle V^{(3)}_{23}\{Q,V^{(3)}_{41}\}\right\rangle={\cal A}(4,2,3,1).&\cr&&({3.25})}$
@@ 3.5. Perturbiner methods
$\eqalignno{{\cal M}_{4}&={1\over 4}\lim_{k_{N}^{2}\rightarrow 0}\left\langle QU_{123}^{(3)}U_{4}^{(3)}\right\rangle,&({3.26})}$
@@
$\eqalignno{G_{\alpha\beta\delta\epsilon}&=E_{(\alpha}{}^{M}E_{\beta}{}^{N}E_{\delta}{}^{P}E_{\epsilon)}{}^{Q}G_{QPNM}&\cr&=H_{\alpha\beta\delta\epsilon}+4\tilde{E}_{(\alpha}{}^{M}\hat{E}_{\beta}{}^{N}\hat{E}_{\delta}{}^{P}\hat{E}_{\epsilon)}{}^{Q}G_{QPNM}+6\tilde{E}_{(\alpha}{}^{M}\tilde{E}_{\beta}{}^{N}\hat{E}_{\delta}{}^{P}\hat{E}_{\epsilon)}{}^{Q}G_{QPNM}&\cr&+4\tilde{E}_{(\alpha}{}^{M}\tilde{E}_{\beta}{}^{N}\tilde{E}_{\delta}{}^{P}\hat{E}_{\epsilon)}{}^{Q}G_{QPNM}+\tilde{E}_{(\alpha}{}^{M}\tilde{E}_{\beta}{}^{N}\tilde{E}_{\delta}{}^{P}\tilde{E}_{\epsilon)}{}^{Q}G_{QPNM}=0,&({3.27})}$
@@
$\eqalignno{\tilde{E}_{A}{}^{M}&=\sum_{{\cal P}}E_{{\cal P},A}{}^{M}e^{k_{\cal{P}}\cdot X},&({3.28})\cr\tilde{E}_{M}{}^{A}&=\sum_{{\cal P}}E_{{\cal P},M}{}^{A}e^{k_{\cal{P}}\cdot X},&({3.29})\cr H_{ABCD}&=\hat{H}_{ABCD}+\sum_{{\cal P}}H_{{\cal{P}},ABCD}e^{k_{\cal{P}}\cdot X},&({3.30})}$
@@
$\eqalignno{H_{\cal{X},\alpha\beta\delta\epsilon}&=4\hat{E}_{\alpha}{}^{M}\hat{E}_{\beta}{}^{N}\hat{E}_{\delta}{}^{P}\hat{E}_{\epsilon}{}^{Q}\hat{E}_{Q}{}^{D}\hat{E}_{P}{}^{C}\hat{E}_{N}{}^{B}E_{{\cal P},M}{}^{A}H_{ABCD}\cr&+6\sum_{{\cal P}U{\cal Q}}\hat{E}_{\alpha}{}^{M}\hat{E}_{\beta}{}^{N}\hat{E}_{\delta}{}^{P}\hat{E}_{\epsilon}{}^{Q}\hat{E}_{Q}{}^{D}\hat{E}_{P}{}^{C}E_{{\cal P},N}{}^{B}E_{{\cal Q},M}{}^{A}H_{ABCD}\cr&+4\sum_{{\cal P}U{\cal Q}U{\cal R}}\hat{E}_{\alpha}{}^{M}\hat{E}_{\beta}{}^{N}\hat{E}_{\delta}{}^{P}\hat{E}_{\epsilon}{}^{Q}\hat{E}_{Q}{}^{D}E_{{\cal P},P}{}^{C}E_{{\cal Q},N}{}^{B}E_{{\cal R},M}{}^{A}H_{ABCD}\cr&+\sum_{{\cal P}U{\cal Q}U{\cal R}U{\cal S}}\hat{E}_{\alpha}{}^{M}\hat{E}_{\beta}{}^{N}\hat{E}_{\delta}{}^{P}\hat{E}_{\epsilon}{}^{Q}E_{{\cal P},Q}{}^{D}E_{{\cal Q},P}{}^{C}E_{{\cal R},N}{}^{B}E_{{\cal S},M}{}^{A}H_{ABCD}.&({3.31})}$
Using the definition $h_{{\cal P},A}{}^{B}=\hat{E}_{A}{}^{M}\tilde{E}_{{\cal P},M}{}^{B}$, one finds the following equation of motion at cubic order in the number of fields
$\eqalignno{H_{\cal{X},\alpha\beta\delta\epsilon}&=-6(\gamma_{ab})_{(\alpha\beta}\sum_{{\cal P}U{\cal Q}}h_{{\cal P},\delta}{}^{a}h_{{\cal Q},\epsilon)}{}^{b}-12(\gamma_{ab})_{(\alpha\kappa}\sum_{{\cal P}U{\cal Q}U{\cal S}}h_{{\cal P},\beta}{}^{\kappa}h_{{\cal Q},\delta}{}^{a}h_{{\cal R},\epsilon)}{}^{b},&({3.32})}$
@@
$\eqalignno{QU_{{\cal X}}^{(3)}&=6(\lambda\gamma^{ab}\lambda)\sum_{{\cal P}U{\cal Q}}\Phi_{{\cal P}}^{a}\Phi_{{\cal Q}}^{b}+12\sum_{{\cal P}U{\cal Q}U{\cal R}}(\lambda\gamma_{ab}\Phi_{{\cal P}})\Phi_{{\cal Q}}^{a}\Phi_{{\cal R}}^{b},&({3.33})}$
where $\Phi_{{\cal P}}^{A}=\lambda^{\alpha}h_{{\cal P},\alpha}{}^{A}$. Note that $\Phi^{A}_{i}=\phi_{i}^{A}$ for any single-particle label $i$.
@@
$\eqalignno{{\cal M}_{4}&=3\bigg{[}\left\langle(\lambda\gamma_{ab}\lambda)\Phi_{12}^{a}\Phi_{3}^{b}U_{4}^{(3)}\right\rangle+\left\langle(\lambda\gamma_{ab}\lambda)\Phi_{13}^{a}\Phi_{2}^{b}U_{4}^{(3)}\right\rangle+\left\langle(\lambda\gamma_{ab}\lambda)\Phi_{23}^{a}\Phi_{1}^{b}U_{4}^{(3)}\right\rangle&\cr&+2\left\langle(\lambda\gamma_{ab}\Phi_{1})\Phi^{a}_{2}\Phi_{3}^{b}U_{4}^{(3)}\right\rangle+2\left\langle(\lambda\gamma_{ab}\Phi_{2})\Phi^{a}_{1}\Phi_{3}^{b}U_{4}^{(3)}\right\rangle+2\left\langle(\lambda\gamma_{ab}\Phi_{3})\Phi^{a}_{1}\Phi_{2}^{b}U_{4}^{(3)}\right\rangle\bigg{]}.&\cr&&({3.34})}$
In order to reproduce the three-particle amplitude (2.47) from the perturbiner formula (3.26) restricted to the case of three external labels, one needs to make the rescaling $\Phi_{ij}^{A}={1\over k_{ij}}\phi^{A}_{ij}$. After doing so, one concludes that ${\cal M}_{4}={\cal A}_{4}$, as promised.

### v2

@@ 3.2. Simplification
$\eqalignno{{\cal A}_{4}^{s}+{\cal A}_{4}^{t}&={6\over s_{12}}\left\langle(\lambda\gamma_{ba}\lambda)\phi^{b}_{12}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+{6\over s_{13}}\left\langle(\lambda\gamma_{ba}\lambda)\phi^{b}_{13}\phi_{2}^{a}U_{4}^{(3)}\right\rangle\cr+&12\left\langle(\lambda\gamma_{ab}\phi_{1})\phi^{a}_{2}\phi_{3}^{b}U_{4}^{(3)}\right\rangle+12\left\langle(\lambda\gamma_{ab}\phi_{2})\phi_{3}^{a}\phi_{1}^{b}U_{4}^{(3)}\right\rangle+12\left\langle(\lambda\gamma_{ab}\phi_{3})\phi_{1}^{a}\phi_{2}^{b}U_{4}^{(3)}\right\rangle\cr+&6\left\langle(\lambda\gamma_{ab}\lambda)\phi_{1}^{a}h^{b}_{(2,c}\phi_{3)}^{c}U_{4}^{(3)}\right\rangle+6\left\langle(\lambda\gamma_{ab}\lambda)\phi_{1}^{a}h^{b}_{(2,\alpha}\phi_{3)}^{\alpha}U_{4}^{(3)}\right\rangle\cr+&6\left\langle C_{1a}\Big(\phi_{(2}^{c}\partial_{c}\phi^{a}_{3)}+\phi_{(2}^{\delta}D_{\delta}\phi^{a}_{3)}-\Omega_{(2}^{\alpha}h^{a}_{3),\alpha}\Big)U_{4}^{(3)}\right\rangle&\cr-&6\left\langle C_{1\alpha}\Big(\phi_{(2}^{c}\partial_{c}\phi^{\alpha}_{3)}+\phi_{(2}^{\delta}D_{\delta}\phi^{\alpha}_{3)}-\Omega_{(2}^{\beta}h^{\alpha}_{3),\beta}\Big)U_{4}^{(3)}\right\rangle.&({3.20})}$
@@
$\eqalignno{{\cal A}^{u}_{4}=&\left\langle\left\{U_{1}^{(3)},V_{23}^{(1)}\right\}U_{4}^{(3)}\right\rangle\cr=&\ {6\over s_{23}}\Big(\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}\tilde{\phi}_{23}^{b}U_{4}^{(3)}\right\rangle+3\left\langle C_{1,a}(\{Q,\tilde{\phi}_{23}^{a}\}-(\lambda\gamma^{a}\tilde{\phi}_{23}))U_{4}^{(3)}\right\rangle\cr&-3\left\langle C_{1,\alpha}([Q,\tilde{\phi}_{23}^{\alpha}]+\tilde{\Omega}^{\alpha}_{23})U_{4}^{(3)}\right\rangle\Big)\cr=&\ {6\over s_{23}}\left\langle(\lambda\gamma_{ab}\lambda)\phi_{23}^{a}\phi^{b}_{1}U_{4}^{(3)}\right\rangle-6\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}h^{b}_{(2,\alpha}\phi^{\alpha}_{3)}U_{4}^{(3)}\right\rangle-6\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}h^{b}_{(2,c}\phi^{c}_{3)}U_{4}^{(3)}\right\rangle\cr&-6\left\langle C_{1,a}\big(\phi_{(2}^{c}\partial_{c}\phi_{3)}^{a}+\phi_{(2}^{\delta}D_{\delta}\phi_{3)}^{a}-\Omega_{(2}^{\alpha}h^{a}_{3),\alpha}\big)U_{4}^{(3)}\right\rangle&\cr&+6\left\langle C_{1,\alpha}\big(\phi_{(2}^{c}\partial_{c}\phi_{3)}^{\alpha}+\phi_{(2}^{\delta}D_{\delta}\phi_{3)}^{\alpha}-\Omega_{(2}^{\beta}h^{\alpha}_{3),\beta}\big)U_{4}^{(3)}\right\rangle.&({3.21})}$
After plugging eqns. (3.20) and (3.21) into eqn. (3.16), one finds the compact formula for the 4-point function in 11D
$\eqalignno{{\cal A}_{4}=&{\cal A}_{4}^{s}+{\cal A}_{4}^{t}+{\cal A}_{4}^{u}&\cr=&\ 6\Bigg[\ {1\over s_{12}}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{12}\phi_{3}^{b}U_{4}^{(3)}\right\rangle+{1\over s_{13}}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{13}\phi_{2}^{b}U_{4}^{(3)}\right\rangle+{1\over s_{23}}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{23}\phi_{1}^{b}U_{4}^{(3)}\right\rangle\cr&\ \ \ \ \ \ \ +2\left\langle(\lambda\gamma_{ab}\phi_{1})\phi^{a}_{2}\phi_{3}^{b}U_{4}^{(3)}\right\rangle+2\left\langle(\lambda\gamma_{ab}\phi_{3})\phi^{a}_{1}\phi_{2}^{b}U_{4}^{(3)}\right\rangle+2\left\langle(\lambda\gamma_{ab}\phi_{2})\phi^{a}_{3}\phi_{1}^{b}U_{4}^{(3)}\right\rangle\Bigg].&\cr&&({3.22})}$
@@ 3.4. BRST invariance and full permutation symmetry
$\eqalignno{&{\cal A}_{4}=4\left\langle V^{(3)}_{12}\{Q,V^{(3)}_{34}\}\right\rangle+4\left\langle V^{(3)}_{13}\{Q,V^{(3)}_{24}\}\right\rangle+4\left\langle V^{(3)}_{23}\{Q,V^{(3)}_{14}\}\right\rangle,&({3.23})}$
@@
$\eqalignno{\{Q,A_{4}\}=&12\left\langle(\lambda\gamma^{[ab}\lambda)(\lambda\gamma^{c]d}\lambda)\phi_{1,a}\phi_{2,b}\phi_{3,c}\phi_{4,d}\right\rangle.&({3.24})}$
@@
$\eqalignno{{\cal A}(1,2,3,4)=&4\left\langle V^{(3)}_{12}\{Q,V^{(3)}_{34}\}\right\rangle+4\left\langle V^{(3)}_{13}\{Q,V^{(3)}_{24}\}\right\rangle+4\left\langle V^{(3)}_{23}\{Q,V^{(3)}_{14}\}\right\rangle\cr=&4\left\langle\{Q,V^{(3)}_{21}\}V^{(3)}_{34}\right\rangle+4\left\langle V^{(3)}_{23}\{Q,V^{(3)}_{14}\}\right\rangle+4\left\langle V^{(3)}_{13}\{Q,V^{(3)}_{24}\}\right\rangle={\cal A}(2,1,3,4),&\cr&&({3.25})}$
@@
$\eqalignno{{\cal A}(1,2,3,4)=&4\left\langle V^{(3)}_{12}\{Q,V^{(3)}_{34}\}\right\rangle+4\left\langle V^{(3)}_{13}\{Q,V^{(3)}_{24}\}\right\rangle+4\left\langle V^{(3)}_{23}\{Q,V^{(3)}_{14}\}\right\rangle\cr=&4\left\langle\{Q,V^{(3)}_{21}\}V^{(3)}_{43}\right\rangle+4\left\langle\{Q,V^{(3)}_{31}\}V^{(3)}_{42}\right\rangle+4\left\langle V^{(3)}_{23}\{Q,V^{(3)}_{41}\}\right\rangle={\cal A}(4,2,3,1).&\cr&&({3.26})}$
@@ 3.5. Perturbiner methods
$\eqalignno{{\cal M}_{4}&={1\over 2}\lim_{k_{4}^{2}\rightarrow 0}\left\langle QU_{123}^{(3)}U_{4}^{(3)}\right\rangle,&({3.27})}$
@@
$\eqalignno{G_{\alpha\beta\delta\epsilon}&=E_{(\alpha}{}^{M}E_{\beta}{}^{N}E_{\delta}{}^{P}E_{\epsilon)}{}^{Q}G_{QPNM}&\cr&=H_{\alpha\beta\delta\epsilon}+4\tilde{E}_{(\alpha}{}^{M}\hat{E}_{\beta}{}^{N}\hat{E}_{\delta}{}^{P}\hat{E}_{\epsilon)}{}^{Q}G_{QPNM}+6\tilde{E}_{(\alpha}{}^{M}\tilde{E}_{\beta}{}^{N}\hat{E}_{\delta}{}^{P}\hat{E}_{\epsilon)}{}^{Q}G_{QPNM}&\cr&+4\tilde{E}_{(\alpha}{}^{M}\tilde{E}_{\beta}{}^{N}\tilde{E}_{\delta}{}^{P}\hat{E}_{\epsilon)}{}^{Q}G_{QPNM}+\tilde{E}_{(\alpha}{}^{M}\tilde{E}_{\beta}{}^{N}\tilde{E}_{\delta}{}^{P}\tilde{E}_{\epsilon)}{}^{Q}G_{QPNM}=0,&({3.28})}$
@@
$\eqalignno{\tilde{E}_{A}{}^{M}&=\sum_{{\cal P}}E_{{\cal P},A}{}^{M}e^{k_{\cal{P}}\cdot X},&({3.29})\cr\tilde{E}_{M}{}^{A}&=\sum_{{\cal P}}E_{{\cal P},M}{}^{A}e^{k_{\cal{P}}\cdot X},&({3.30})\cr H_{ABCD}&=\hat{H}_{ABCD}+\sum_{{\cal P}}H_{{\cal{P}},ABCD}e^{k_{\cal{P}}\cdot X},&({3.31})}$
@@
$\eqalignno{H_{\cal{X},\alpha\beta\delta\epsilon}&=4\hat{E}_{\alpha}{}^{M}\hat{E}_{\beta}{}^{N}\hat{E}_{\delta}{}^{P}\hat{E}_{\epsilon}{}^{Q}\hat{E}_{Q}{}^{D}\hat{E}_{P}{}^{C}\hat{E}_{N}{}^{B}E_{{\cal X},M}{}^{A}H_{ABCD}\cr&+6\sum_{{\cal P}\cup{\cal Q}={\cal X}}\hat{E}_{\alpha}{}^{M}\hat{E}_{\beta}{}^{N}\hat{E}_{\delta}{}^{P}\hat{E}_{\epsilon}{}^{Q}\hat{E}_{Q}{}^{D}\hat{E}_{P}{}^{C}E_{{\cal P},N}{}^{B}E_{{\cal Q},M}{}^{A}H_{ABCD}\cr&+4\sum_{{\cal P}\cup{\cal Q}\cup{\cal R}={\cal X}}\hat{E}_{\alpha}{}^{M}\hat{E}_{\beta}{}^{N}\hat{E}_{\delta}{}^{P}\hat{E}_{\epsilon}{}^{Q}\hat{E}_{Q}{}^{D}E_{{\cal P},P}{}^{C}E_{{\cal Q},N}{}^{B}E_{{\cal R},M}{}^{A}H_{ABCD}\cr&+\sum_{{\cal P}\cup{\cal Q}\cup{\cal R}\cup{\cal S}={\cal X}}\hat{E}_{\alpha}{}^{M}\hat{E}_{\beta}{}^{N}\hat{E}_{\delta}{}^{P}\hat{E}_{\epsilon}{}^{Q}E_{{\cal P},Q}{}^{D}E_{{\cal Q},P}{}^{C}E_{{\cal R},N}{}^{B}E_{{\cal S},M}{}^{A}H_{ABCD}.&({3.32})}$
Using the definition ${\tenit h}_{{\cal P},A}{}^{B}=\hat{E}_{A}{}^{M}\tilde{E}_{{\cal P},M}{}^{B}$, one finds the following equation of motion at cubic order in the number of fields
$\eqalignno{H_{\cal{X},\alpha\beta\delta\epsilon}&=-6(\gamma_{ab})_{(\alpha\beta}\sum_{{\cal P}\cup{\cal Q}={\cal X}}h_{{\cal P},\delta}{}^{a}h_{{\cal Q},\epsilon)}{}^{b}-12(\gamma_{ab})_{(\alpha\kappa}\sum_{{\cal P}\cup{\cal Q}\cup{\cal S}={\cal X}}h_{{\cal P},\beta}{}^{\kappa}h_{{\cal Q},\delta}{}^{a}h_{{\cal R},\epsilon)}{}^{b},&\cr&&({3.33})}$
@@
$\eqalignno{QU_{{\cal X}}^{(3)}&=6(\lambda\gamma^{ab}\lambda)\sum_{{\cal P}\cup{\cal Q}={\cal X}}{\tenit\Phi}_{{\cal P}}^{a}{\tenit\Phi}_{{\cal Q}}^{b}+12\sum_{{\cal P}\cup{\cal Q}\cup{\cal R}={\cal X}}(\lambda\gamma_{ab}{\tenit\Phi}_{{\cal P}}){\tenit\Phi}_{{\cal Q}}^{a}{\tenit\Phi}_{{\cal R}}^{b},&({3.34})}$
where ${\tenit\Phi}_{{\cal P}}^{A}=\lambda^{\alpha}{\tenit h}_{{\cal P},\alpha}{}^{A}$. Note that ${\tenit\Phi}^{A}_{i}=\phi_{i}^{A}$ for any single-particle label $i$.
@@
$\eqalignno{{\cal M}_{4}&=6\bigg[\left\langle(\lambda\gamma_{ab}\lambda){\tenit\Phi}_{12}^{a}{\tenit\Phi}_{3}^{b}U_{4}^{(3)}\right\rangle+\left\langle(\lambda\gamma_{ab}\lambda){\tenit\Phi}_{13}^{a}{\tenit\Phi}_{2}^{b}U_{4}^{(3)}\right\rangle+\left\langle(\lambda\gamma_{ab}\lambda){\tenit\Phi}_{23}^{a}{\tenit\Phi}_{1}^{b}U_{4}^{(3)}\right\rangle&\cr&+2\left\langle(\lambda\gamma_{ab}{\tenit\Phi}_{1}){\tenit\Phi}^{a}_{2}{\tenit\Phi}_{3}^{b}U_{4}^{(3)}\right\rangle+2\left\langle(\lambda\gamma_{ab}{\tenit\Phi}_{2}){\tenit\Phi}^{a}_{1}{\tenit\Phi}_{3}^{b}U_{4}^{(3)}\right\rangle+2\left\langle(\lambda\gamma_{ab}{\tenit\Phi}_{3}){\tenit\Phi}^{a}_{1}{\tenit\Phi}_{2}^{b}U_{4}^{(3)}\right\rangle\bigg].&\cr&&({3.35})}$
In order to reproduce the four-particle amplitude (3.22) from the perturbiner formula (3.27) restricted to the case of four external labels, one needs to make the rescaling ${\tenit\Phi}_{ij}^{A}={1\over s_{ij}}\phi^{A}_{ij}$, which is consistent with our discussion in Appendix B. After doing so, one concludes that ${\cal M}_{4}={\cal A}_{4}$, as promised.